### v1

# Sharp concentration inequality for the sum of random variables

Abstract
We present a universal concentration bound for sums of random variables under arbitrary dependence, and we prove it is asymptotically optimal for every fixed common marginal law. The concentration bound is a direct—yet previously unnoticed—consequence of the subadditivity of expected shortfall, a property well known to financial statisticians. The sharpness result is a significant contribution relying on the construction of worst-case dependency profiles between identically distributed random variables.
1 Maximally nonincreasing operator framework and main results
Following the approach of Rockafellar and Royset (2014) on superquantiles and our previous work on concentration inequalities Louart (2024), we encode concentration statements as inequalities between maximally nonincreasing operators rather than as pointwise inequalities between real-valued functions. The operator viewpoint is convenient here because objects underlying tail bounds (like survival functions and quantiles) are monotone but may have jumps (like the limiting profile of the law of large numbers); representing them as set-valued operators provides a canonical way to handle discontinuities without choosing arbitrary versions (lower or upper semicontinuous).
Concretely, we work in the class of “maximally nonincreasing operators” generalizing nonincreasing mappings on $\mathbb{R}$, that we denote $\mathcal{M}_{\downarrow}$ and that contains set-valued mappings $\alpha:\mathbb{R}\to 2^{\mathbb{R}}$ satisfying [^1] The domain, range, and graph of an operator $\alpha:\mathbb{R}\to 2^{\mathbb{R}}$ are respectively defined as $\operatorname{Dom}(\alpha)=\{t\in\mathbb{R}\ |\ \alpha(t)\neq\emptyset\}$, $\operatorname{Ran}(\alpha)=\cup_{t\in\mathbb{R}}\alpha(t)$, and $\operatorname{Gra}(\alpha)=\{(x,y)\,:\,x\in\operatorname{Dom}(\alpha),\ y\in\alpha(x)\}$. The domain and range of a maximally nonincreasing operator are intervals; moreover, for any $t\in\mathbb{R}$, $\alpha(t)$ is a closed interval.:
$\forall(x,u),(y,v)\in\operatorname{Gra}(\alpha):\quad(y-x)(u-v)\geq 0\qquad\text{and}\qquad\operatorname{Dom}(\alpha)-\operatorname{Ran}(\alpha)=\mathbb{R}.$
Further characterizations can be found in Bauschke and Combettes (2011); Louart (2024).
In what follows, $\mu$ designates a probability measure on $\mathbb{R}$ and $X\sim\mu$, a given random variable. Operator inversion allows one to move naturally between the survival operator $S_{X}:\mathbb{R}\to 2^{\mathbb{R}}$ (also denoted $S_{\mu}$) and the tail quantile operator $T_{X}:=S_{X}^{-1}:\mathbb{R}\to 2^{\mathbb{R}}$ (also denoted $T_{\mu}$), both maximally nonincreasing and defined respectively by
$\forall t\in\mathbb{R}:\qquad$ $S_{X}(t):=[\mathbb{P}(X>t),\mathbb{P}(X\geq t)]=[\mu(]t,+\infty[),\mu(]t,+\infty[)]=:S_{\mu}(t)$
$\forall p\in[0,1]:\qquad$ $T_{X}(p):=\{t\in\mathbb{R}\ :\ p\in S_{X}(t)\}=:T_{\mu}(p).$
The natural order relation to set concentration inequalities on survival operator was introduced in Louart (2024):
Definition 1 (Interval Order and Point-wise resolvent Order between operators).
The order between intervals $A,B\subset\mathbb{R}$ is defined by
$A\leq B\quad\Longleftrightarrow\quad B_{+}\subset A_{+}\ \ \text{and}\ \ A_{-}\subset B_{-},$
with $A_{+}=\{x\in\mathbb{R}\ :\ \exists y\in A,\ y\leq x\}$ and $A_{-}=\{x\in\mathbb{R}\ :\ \exists y\in A,\ y\geq x\}$.
Given $f,g\in\mathcal{M}_{\downarrow}$, we write $f\leq g$ if and only if
$\operatorname{Dom}(f)\leq\operatorname{Dom}(g)\qquad\text{and}\qquad\forall x\in\operatorname{Dom}(f)\cap\operatorname{Dom}(g):\ f(x)\leq g(x),$
where $\operatorname{Dom}(f)$, $\operatorname{Dom}(g)$, $f(x)$ and $g(x)$ are (by construction) intervals (see (Bauschke and Combettes, 2011, Proposition 20.31, Corollary 21.12)).
A confortable advantage of working with survival and tail quantile operators instead of working with cumulative distribution operator $t\mapsto 1-S_{X}(t)$ and quantile operator $p\mapsto T_{X}(1-p)$ as in Rockafellar and Royset (2014) is that an upper bound on the survival profile is equivalent to the corresponding upper bound on the tail quantile profile. More precisely, for any $\alpha\in\mathcal{M}_{\downarrow}$,
$S_{X}\leq\alpha\qquad\Longleftrightarrow\qquad T_{X}\leq\alpha^{-1}.$ (1)
Our major task then expresses:
  • Task:
    Given $n$ probability measures $\mu_{1},\ldots,\mu_{n}$ on $\mathbb{R}$, construct a minimal operator $\beta_{\mu_{1},\ldots,\mu_{n}}\in\mathcal{M}_{\downarrow}$ such that, for any collection of $n$ random variables $X_{1}\sim\mu_{1},\ldots,X_{n}\sim\mu_{n}$:
    $S_{X_{1}+\cdots+X_{n}}\leq\beta_{\mu_{1},\ldots,\mu_{n}}.$
Assuming $\mu_{1}=\cdots=\mu_{n}$ and (i) considering $X_{1},\ldots,X_{n}$ i.i.d, (ii) considering $X_{1}=\cdots=X_{n}$ and (iii) using the union bound, on gets the first simple bound [^2] Maximum of operators is given a rigorous definition in Louart (2024) that is not necessary to provide here since this inequality just plays a heuristic role.
$\max(\mathrm{Incr}_{\mathbb{E}[X]},S_{X})\leq\beta_{\mu,\ldots,\mu}\leq nS_{X},$
where $X\sim\mathbb{\mu}$ and for any $\delta\in\mathbb{R}$, we denote by $\mathrm{Incr}_{\delta}\in\mathcal{M}_{\downarrow}$ the operator defined by
$\mathrm{Incr}_{\delta}(t)=\begin{cases}\{1\},&t<\delta,\\ [0,1],&t=\delta,\\ \{0\},&t>\delta.\end{cases}$
[…]
$\forall p\in\mathring{\operatorname{Dom}}(f)=\mathring{\operatorname{Dom}}(\mathcal{H}(f)):\qquad\mathcal{H}(f)(p)=\frac{1}{p}\int_{0}^{p}f(r)\,dr=\int_{0}^{1}f(pr)\,dr,$
then, since $\mathcal{H}(f)\in\mathcal{M}_{\downarrow}$, it can be shown that:
  • if $0\in\operatorname{Dom}(f)$, then $0\in\operatorname{Dom}(\mathcal{H}(f))$ and $\mathcal{H}(f)(0)=f(0)$,
  • denoting $p_{f}:=\sup(\operatorname{Dom}(f))$, we have $p_{f}\in\operatorname{Dom}(\mathcal{H}(f))$ and $\mathcal{H}(f)(p_{f})=\left]-\infty,\frac{1}{p_{f}}\int_{0}^{p_{f}}f(r)\,dr\right]$.
An example of a Hardy transform is depicted on Figure 1. We record a few properties that will be useful later:
[…]
    $0,1]\subset\operatorname{Dom}(\mathcal{H}(f))$ and for all $t\in]0,1[$, $\mathcal{H}(f)(t)$ is a singleton;
[…]
    $\mathcal{H}$ is linear;
[…]
    $f\leq\mathcal{H}(f)$;
[…]
    if $f$ is convex then $\mathcal{H}(f)$ is convex;
[…]
    if $f\leq g$ then $\mathcal{H}(f)\leq\mathcal{H}(g)$.
A first partial answer to our Task is then given by
Theorem 1.
Given $n$ random variables $X_{1},\ldots,X_{n}$ admitting expectations, one can bound:
$S_{X_{1}+\cdots+X_{n}}\leq\left(\mathcal{H}(T_{X_{1}})+\cdots+\mathcal{H}(T_{X_{n}})\right)^{-1}.$
As a comparison, in (Louart, 2024, Proposition 42) we only proved the weaker bound (a generalization of union bound to the non-identically distributed case):
$S_{X_{1}+\cdots+X_{n}}\leq n\bigl(T_{X_{1}}+\cdots+T_{X_{n}}\bigr)^{-1}.$
When $X_{1},\ldots,X_{n}\sim\mu$ are identically distributed, Theorem 1 rewrites:
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq\mathcal{H}(T_{\mu})^{-1}$ (2)
as expected, removing completely the dependence on $n$ of the bound.
Remark 1.
It is instructive to understand what is lost at the level of moments when applying Theorem 1. For any nonnegative random variable $X\sim\mu$, for some probability measure $\mu$ on $\mathbb{R}$ one has the identity (see for instance (Rockafellar and Royset, 2014, (3.4))):
$M_{q}(\mu):=\mathbb{E}[X^{q}]=\int_{0}^{\infty}\mathbb{P}(X^{q}\geq t)\,dt=\int_{0}^{\infty}T_{X}(p)^{q}\,dp.$ (3)
Following Hardy’s inequality (Hardy et al., 1952, Theorem 9.8.2), valid for any measurable function $f:]0,1[\to\mathbb{R}_{+}$,
$\int_{0}^{1}\mathcal{H}(f)(t)^{q}\,dt\leq\left(\frac{q}{q-1}\right)^{q}\int_{0}^{1}f(t)^{q}\,dt.$
Given an operator $\beta\in\mathcal{M}_{\downarrow}$ with $\operatorname{Dom}(\beta)\subset\mathbb{R}_{+}$, we define the “moments” of $\beta$ by
$M_{q}(\beta):=\int_{\mathbb{R}_{+}}\beta\!\left(t^{\frac{1}{q}}\right)\,dt,\quad q>0,$ (4)
so that $M_{q}(S_{X})=\mathbb{E}[X^{q}]=M_{q}(\mu)$. Then, given $n$ random variables $X_{1},\ldots,X_{n}\sim\mu$, Theorem 1 applied in the $(X_{i})_{i\in[k]}$-identically distributed setting (2) yields the moment bound
$M_{q}\left(S_{\frac{1}{n}\sum_{i=1}^{n}X_{k}}\right)$ $=\int_{0}^{1}T_{\frac{1}{n}\sum_{i=1}^{n}X_{k}}(p)^{q}\,dp\leq\int_{0}^{1}\mathcal{H}(T_{X})(p)^{q}\,dp$
$\leq\left(\frac{q}{q-1}\right)^{q}\int_{0}^{1}T_{X}(p)^{q}\,dp=\left(\frac{q}{q-1}\right)^{q}M_{q}(S_{X})$ (5)
This inequality is clearly underperforming since a mere application of Jensen inequality already yields for all $q>1$ the moment optimal bound [^6] It is reached for instance in the case $X_{1}=\cdots=X_{n}$.:
$M_{q}(S_{\frac{1}{n}\sum_{i=1}^{n}X_{k}})\leq M_{q}(S_{X})$
Still (1) demonstrates that very little is lost through application of Theorem 1. Anyway, the strength of Theorem 1 is that it provides a universal tail bound, valid for all thresholds $t$ and all joint distributions of $(X_{1},\ldots,X_{n})$ with given marginals; for this reason, moments are not the most relevant metric here.
To pursue the study of the scope of Theorem 1, we give below a general result that efficiently bounds the concentration of the sum for a wide range of distributions. We employ the shorthand notation (already introduced in Louart (2024)) $\mathrm{Id}^{-a}:\mathbb{R}\to 2^{\mathbb{R}}$ defined by
$\forall t>0:\quad\mathrm{Id}^{-a}(t)=\{t^{-a}\},\qquad\text{and}\qquad\mathrm{Id}^{-a}((-\infty,0])=\emptyset.$
Corollary 1.
Given a probability measure $\mu$ on $\mathbb{R}$ and $n$ identically distributed random variables $X_{1},\ldots,X_{n}\sim\mu$, assume that there exist $\alpha\in\mathcal{M}_{\downarrow}$ satisfying $\operatorname{Ran}(\alpha)\subset\mathbb{R}^{+}_{*}$ and $q>1$ such that [^7] We say that an operator $g:\mathbb{R}\to 2^{\mathbb{R}}$ is convex if any real-valued mapping $h:\operatorname{Dom}(g)\to\mathbb{R}$ satisfying $\forall t\in\operatorname{Dom}(g)$, $h(t)\in g(t)$ is convex.:
$S_{\mu}\leq\alpha\qquad\text{and}\qquad\mathrm{Id}^{-\frac{1}{q}}\circ\alpha\ \ \text{is convex}.$
Then one can bound
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq\left(\frac{q}{q-1}\right)^{q}\alpha.$
If we assume in addition [^8] The second result of Theorem 3 in the next section states that $\log\circ\alpha$ convex implies $\mathrm{Id}^{-\frac{1}{q}}\circ\alpha$ convex for all $q>0$. that $\log\circ\alpha$ is convex, then
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq e\,\alpha.$
This corollary is proved in Section 3.
Theorem 1 is, in fact, a reformulation of a property that is classical in financial mathematics: the subadditivity of expected shortfall Acerbi and Tasche (2002) (also called conditional value-at-risk Rockafellar and Uryasev (2002), and more recently superquantiles Rockafellar and Royset (2014)). Subadditivity is one of the coherence axioms introduced in Artzner et al. (1999) for risk measures. This property was obtained simultaneously in (Acerbi and Tasche, 2002, Proposition 3.1) (under the name “shortfall expectation”) and in (Rockafellar and Uryasev, 2002, Corollary 12) (under the name “conditional value-at-risk”). Although these notions were initially presented differently, they were later shown to coincide; in our notation, they correspond to the mapping $p\mapsto\mathcal{H}(T_{X})(1-p)$. Subadditivity can be written as
$\mathcal{H}(T_{X_{1}+\cdots+X_{n}})\leq\mathcal{H}(T_{X_{1}})+\cdots+\mathcal{H}(T_{X_{n}}).$
Noting that $T_{X_{1}+\cdots+X_{n}}\leq\mathcal{H}(T_{X_{1}+\cdots+X_{n}})$ and relying on (1), we recover exactly Theorem 1.
Up to this point, no structurally new phenomenon has been introduced; we now have all the ingredients to state our main contribution: asymptotic sharpness of the inequality provided by Theorem 1. Before stating the result, if $\mathbb{E}[|X|]<\infty$, let us introduce the quantity
$\forall p\in]0,1[:\qquad\Delta_{\mu}(p):=\frac{\mathcal{H}(T_{\mu})(p)-\mathbb{E}[X]}{1-p}.$ (6)
Lemma 1.
Given any probability measure $\mu$ on $\mathbb{R}$ with $\mathbb{E}[|X|]<\infty$, any $X\sim\mu$, and any $p\in]0,1[$,
$\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p)\leq\mathbb{E}[X]\leq\mathcal{H}(T_{\mu})(p),$ (7)
with equality in (7) if and only if $\Delta_{\mu}(p)=0$.
Proof.
First note that $\Delta_{\mu}(p)\geq 0$ since $\mathcal{H}(T_{\mu})$ is nonincreasing and $\mathcal{H}(T_{\mu})(p)\geq\mathcal{H}(T_{\mu})(1)=\mathbb{E}[X]$. Moreover,
$\mathbb{E}[X]-\bigl(\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p)\bigr)=\frac{p}{1-p}\,\bigl(\mathcal{H}(T_{\mu})(p)-\mathbb{E}[X]\bigr)\geq 0,$
which proves (7) and the characterization of the equality case. ∎
Define the limiting survival operator $S_{\mu,p}\in\mathcal{M}_{\downarrow}$ satisfying on continuity points:
$S_{\mu,p}(t)=\begin{cases}\{1\},&\text{if }t<H(T_{\mu})(p)-\Delta_{\mu}(p),\\ \{p\},&\text{if }H(T_{\mu})(p)-\Delta_{\mu}(p)<t<H(T_{\mu})(p),\\ \{0\},&\text{if }t>H(T_{\mu})(p).\end{cases}$
Finally, we set $S_{\mu,1}:=\mathrm{Incr}_{\mathbb{E}[X]}$ (note that $\mathcal{H}(T_{\mu})(1)=\mathbb{E}[X]$). In particular, the sharpness will be shown at the contact point:
$p\in S_{\mu,p}(H(T_{\mu})(p)),$
that we represented by a green dot on the tail-quantile profiles graph provided on Figure 1
Figure 1: Representation of $T_{X}$, $\mathrm{Incr}_{\mathbb{E}[X]}$, $\mathcal{H}(T_{X})$ and $S_{\mu,p}^{-1}$ and the asymptotic contact point.
A final answer to our Task (in the identically distributed and finite expectation case) is given by
Theorem 2.
Given a probability distribution $\mu$ on $\mathbb{R}$ admitting a finite expectation, and given any $p\in(0,1]$, there exists a sequence of identically distributed random variables $(X_{n})_{n\in\mathbb{N}}\in\mathbb{R}^{\mathbb{N}}$ such that for all $n\in\mathbb{N}$, $X_{n}\sim\mu$ and [^9] That is, for any continuity point of $S_{\mu,p}$. for all $t\in\mathbb{R}\setminus\{a_{\mu}(p),b_{\mu}(p)\}$,
$S_{\frac{1}{n}(X_{1}+\cdots+X_{n})}(t)\underset{n\to\infty}{\longrightarrow}S_{\mu,p}(t).$
This theorem will be proved constructively in Section 2.
The question remains open as to whether our sum concentration bound can be attained for finite $n$. One should presumably restrict to continuous distributions, since counterexamples are easy to find when $\mu$ has atoms. Moreover, when $n=2$, if we want to attain $S_{\mu,1}=\mathrm{Incr}_{\mathbb{E}[X]}$ with $S_{\frac{X_{1}+X_{2}}{2}}$ for $X_{1},X_{2}\sim\mu$, we must have $X_{1}=2\mathbb{E}[X]-X_{2}$ almost surely. Thus the concentration inequality can only be reached for $\mu$ symmetric around its expectation (and this is only for the threshold $p=1$). More freedom is available for $n\geq 3$, but the problem appears difficult and, having no specific application in mind, we do not explore this direction.
2 Mixed slot variables to sharpnen sum concentration tail
Given $n\in\mathbb{N}$, $p\in]0,1[$ and a probability measure $\mu$ on $\mathbb{R}$, introduce the following notations to prepare the construction [^10] The choice of $u_{0}=u_{0}^{\prime}$ is not important because $\mu(]\inf T_{\mu}(p),\sup T_{\mu}(p)[)=0$.. Figure 2 gives a graphical representation of this choice:
  • $p_{0}=p_{0}^{\prime}:=p$ and, for all $i\in[n]$,
    $p_{i}:=\frac{n-i}{n}p,\qquad p_{i}^{\prime}:=\frac{i}{n}+\frac{n-i}{n}p,$
  • for all $i\in[n]$, pick any
    $u_{i}\in T_{X}(p_{i})$ $\operatorname{and}$ $u_{i}^{\prime}\in T_{X}(p_{i}^{\prime}),$
    if $T_{X}(1)=\emptyset$ then take $u_{n}^{\prime}=$-$\infty$, and if $T_{X}(0)=\emptyset$, take $u_{n}=+\infty$.
  • pick $u_{0}=u_{0}^{\prime}\in T_{X}(p)$,
Figure 2: (Left) detail of the choice of of the different parameters for a given tail quantile operator and (Right) cyclic mixing of slot variables: value of $X_{1},\ldots,X_{n},\sum_{k=1}^{n}X_{k}$ depending on the values of $\epsilon_{1},\ldots,\epsilon_{n},\epsilon_{1}^{\prime},\ldots,\epsilon_{n}^{\prime}$.
Note that:
  • (P1a)
    $\infty\leq u_{n}^{\prime}\leq u_{n-1}^{\prime}\leq\dots\leq u_{0}^{\prime}=u_{0}\leq\dots\leq u_{n-1}\leq u_{n}\leq+\infty$,
  • (P1b)
    The set $\bigcup_{i\in[n]}[u_{i}^{\prime},u_{i-1}^{\prime}[\ \cup\ \bigcup_{i\in[n]}[u_{i-1},u_{i}[$ forms a partition of $\mathbb{R}$,
For any $i\in[n]$, we construct two probability measures $\mu_{i},\mu_{i}^{\prime}$ on $\mathbb{R}$ through their survival function, respectively by setting, forall $t\in\mathbb{R}$:
$S_{\mu_{i}}(t)=\begin{cases}1&\text{if }t<u_{i-1},\\ \frac{S_{\mu}(t)-p_{i}}{p_{i-1}-p_{i}}&\text{if }u_{i-1}<t<u_{i},\\ 0&\text{if }t>u_{i}\end{cases}$ $\operatorname{and}$ $S_{\mu_{i}^{\prime}}(t)=\begin{cases}1&\text{if }t<u_{i}^{\prime},\\ \frac{S_{\mu}(t)-p_{i-1}^{\prime}}{p_{i}^{\prime}-p_{i-1}^{\prime\prime}}&\text{if }u_{i}^{\prime}<t<u_{i-1}^{\prime},\\ 0&\text{if }t>u_{i-1}^{\prime},\end{cases}$ (8)
where the intermediate set value denotes, in both cases, an affine image of the interval $S_{\mu}(t)$. One can check that with this definitions:
  • (P2a)
    if $A\cap[u_{i-1},u_{i}]=\emptyset$ (resp. $A\cap[u_{i}^{\prime},u_{i-1}^{\prime}]=\emptyset$), then $\mu_{i}(A)=0$ (resp. $\mu_{i}^{\prime}(A)=0$),
  • (P2b)
    if $u_{i}=u_{i-1}$ (resp. $u_{i}^{\prime}=u_{i-1}^{\prime}$), then $\mu_{i}(A)=\mathbbm{1}(u_{i}\in A)$ (resp. $\mu_{i}^{\prime}(A)=\mathbbm{1}(u_{i}^{\prime}\in A)$).
As defined, $\mu_{i}$ and $\mu_{i}^{\prime}$ can be viewed as “slot restrictions” of $\mu$, as stated by the next lemma.
Lemma 2.
Given $i\in[n]$ and $s\in]0,1[$,
$T_{\mu_{i}}(s)=T_{\mu}\!\left(p_{i}+\frac{p}{n}s\right),\qquad T_{\mu_{i}^{\prime}}(s)=T_{\mu}\!\left(p_{i-1}^{\prime}+\frac{1-p}{n}s\right).$
Proof.
We prove the first identity (the second one is analogous). If $u_{i}=u_{i-1}$, then by (P2b) the measure $\mu_{i}$ is the Dirac mass at $u_{i}$, hence $T_{\mu_{i}}(s)=\{u_{i}\}$ for all $s\in]0,1[$. Since $T_{\mu}$ is constant on $(p_{i},p_{i-1})$ in that case, the identity follows.
Assume now that $u_{i}\neq u_{i-1}$. Following the formulas given in (8), for any $s\in]0,1[$ and $t\in\mathbb{R}$,
$s\in S_{\mu_{i}}(t)\quad\Longleftrightarrow\quad p_{i}+(p_{i-1}-p_{i})s\in S_{\mu}(t)\quad\Longleftrightarrow\quad t\in T_{\mu}\!\left(p_{i}+\frac{p}{n}s\right),$
since $p_{i-1}-p_{i}=\frac{p}{n}$. This proves $T_{\mu_{i}}(s)=T_{\mu}\!\left(p_{i}+\frac{p}{n}s\right)$. ∎
Define now “independent slot variables” $W_{n},W_{n-1},\dots,W_{1},W^{\prime}_{1},\dots,W^{\prime}_{n-1},W^{\prime}_{n}$ such that for any $i\in[n]$:
$S_{W_{i}}=S_{\mu_{i}}\qquad\operatorname{and}\qquad S_{W_{i}^{\prime}}=S_{\mu_{i}^{\prime}}.$
We call them “slot variables” since, by (P2a) :
$W_{i}\in[u_{i-1},u_{i}]$ $\operatorname{and}$ $W_{i}^{\prime}\in[u_{i}^{\prime},u_{i-1}^{\prime}]$ $a.s.$ (9)
Lemma 3.
With the notation of (6),
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}W_{i}\right]=\mathcal{H}(T_{\mu})(p)\qquad\text{and}\qquad\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}\right]=\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p).$
Proof.
By Lemma 2 and the identity $\mathbb{E}[Y]=\int_{0}^{1}T_{Y}(s)\,ds$, true for integrable $Y$ (see for instance (Rockafellar and Royset, 2014, (3.4))),
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}W_{i}\right]$ $=\frac{1}{n}\sum_{i=1}^{n}\int_{0}^{1}T_{\mu_{i}}(s)\,ds=\frac{1}{n}\sum_{i=1}^{n}\frac{n}{p}\int_{p_{i}}^{p_{i-1}}T_{\mu}(u)\,du=\frac{1}{p}\int_{0}^{p}T_{\mu}(u)\,du=\mathcal{H}(T_{\mu})(p).$
Similarly,
$\mathbb{E}\left[\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}\right]$ $=\frac{1}{n}\sum_{i=1}^{n}\int_{0}^{1}T_{\mu_{i}^{\prime}}(s)\,ds=\frac{1}{n}\sum_{i=1}^{n}\frac{n}{1-p}\int_{p_{i-1}^{\prime}}^{p_{i}^{\prime}}T_{\mu}(u)\,du$
$=\frac{1}{1-p}\int_{p}^{1}T_{\mu}(u)\,du=\frac{\mathbb{E}[\mu]}{1-p}-\frac{p}{1-p}\mathcal{H}(T_{\mu})(p)=\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p).$
where we recall the notation introduced in (6): $\Delta_{\mu}(p):=\frac{\mathcal{H}(T_{\mu})(p)-\mathbb{E}[X]}{1-p}$. ∎
Actually, we have a law of large number (LLN) for those slot variables
Lemma 4.
If we assume that $\mu$ admits an expectation than, for any $p\in]0,1[$ and any $t>0$:
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{i=1}^{n}W_{i}-\mathcal{H}(T_{\mu})(p)\right|\geq t\right)\underset{n\to\infty}{\longrightarrow}0$
$\mathbb{P}\left(\left|\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}-\mathcal{H}(T_{\mu})(p)+\Delta_{\mu}(p)\right|\geq t\right)\underset{n\to\infty}{\longrightarrow}0$
Proof.
We just prove the first result since the LLN for $\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}$ is proven the same way. Let us first exclude the last elements of the sum that can be unbounded, then we can bound alsmost surely thanks to (9):
$\frac{1}{n}\sum_{i=1}^{n-1}u_{i}\leq\frac{1}{n}\sum_{i=1}^{n-1}W_{i}\leq\frac{1}{n}\sum_{i=1}^{n-1}u_{i-1},$
and by definition of the integral of operators (which is a natural extention of the inntegral of measurable mappings defined as the limit of integral of simple functions as we see them apppear here), we see that:
$\frac{1}{n}\sum_{i=1}^{n-1}W_{i}\underset{n\to\infty}{\longrightarrow}0\qquad a.s$
Besides we know from definition of $S_{\mu_{n}}$ in (8) that ($p_{n}=0$):
$\mathbb{P}\left(\left|\frac{W_{n}}{n}\right|\geq t\right)=S_{\mu_{i}}(tn)\leq\frac{S_{\mu}(tn)}{p_{n-1}}=\frac{n}{p}S_{\mu}(tn)=\frac{n}{p}\int_{tn}^{\infty}d\mu(u)\leq\frac{1}{pt}\int_{tn}^{\infty}ud\mu(u)\underset{n\to\infty}{\longrightarrow}0,$
by continuity of the integral since $u\to u$ is integrable under $\mu$ by assumption. ∎
We now construct $n$ dependent random variables following law $\mu$ by mixing the slot variables. Let $(\epsilon_{1},\ldots,\epsilon_{n},\epsilon_{1}^{\prime},\ldots,\epsilon_{n}^{\prime})\in\{0,1\}^{2n}$ be a random vector, independent of $(W_{i})_{i\in[n]}$ and $(W_{i}^{\prime})_{i\in[n]}$, such that:
  1. 1.
    $\mathbb{P}\!\left(\sum_{i=1}^{n}\epsilon_{i}+\sum_{i=1}^{n}\epsilon_{i}^{\prime}=1\right)=1$,
  2. 2.
    $\mathbb{P}(\epsilon_{i}=1)=\frac{p}{n}$ and $\mathbb{P}(\epsilon_{i}^{\prime}=1)=\frac{1-p}{n}$ for all $i\in[n]$.
For any $k,i\in[n]$, define the cyclic permutation
$\sigma_{k}(i)=\begin{cases}i+k-1,&\text{if }i+k-1\leq n,\\ i+k-n-1,&\text{if }n+1\leq i+k-1\ (\leq 2n-1).\end{cases}$
Finally, following the description given on Figure 2, define
$\forall k\in[n]:\qquad X_{k}=\sum_{i=1}^{n}\epsilon_{\sigma_{k}(i)}W_{i}+\sum_{i=1}^{n}\epsilon_{\sigma_{k}(i)}^{\prime}W_{i}^{\prime}.$ (10)
Lemma 5.
For all $k\in[n]$, $X_{k}\sim\mu$.
Proof.
Since $(\epsilon_{i})_{i}$ and $(\epsilon_{i}^{\prime})_{i}$, $(W_{i})_{i}$ and $(W_{i}^{\prime})_{i}$ are all independent, we can express, for any $t\in\mathbb{R}$:
$S_{X_{k}}(t)=\sum_{i=1}^{n}\mathbb{P}(\epsilon_{\sigma_{k}(i)}=1)\,S_{\mu_{i}}(t)+\sum_{i=1}^{n}\mathbb{P}(\epsilon_{\sigma_{k}(i)}^{\prime}=1)\,S_{\mu_{i}^{\prime}}(t)$
Now if, say, $t\in[u_{i-1},u_{i}]$, we know that
  • $\forall j<i$: $S_{\mu_{j}}(t)=0$,
  • $\forall j\in[n]$: $S_{\mu_{j}^{\prime}}(t)=0$ and $P(\epsilon_{\sigma_{k}(i)})=1$,
  • $\forall j>i$: $S_{\mu_{j}}(t)=1$,
  • $S_{\mu_{i}}(t)=\frac{S_{\mu}(t)-p_{i}}{p_{i-1}-p_{i}}$.
That allows us to evaluate:
$S_{X_{k}}(t)=\frac{p}{n}\frac{S_{\mu}(t)-p_{i}}{p_{i-1}-p_{i}}+\frac{n-i}{n}\frac{p}{n}=S_{\mu}(t),$
thanks to the identity $p_{i-1}-p_{i}=\frac{p}{n}$ and the definition of $p_{i}$. The same thing happens if $t$ belongs to any other slot given in (P1b) . ∎
Proof of Theorem 2.
Using (10) and the fact that $\sum_{k=1}^{n}\epsilon_{\sigma_{k}(i)}=\sum_{j=1}^{n}\epsilon_{j}$ for each fixed $i$, we obtain
$\frac{1}{n}\sum_{k=1}^{n}X_{k}$ $=\frac{1}{n}\left(\sum_{i=1}^{n}\left(\sum_{k=1}^{n}\epsilon_{\sigma_{k}(i)}\right)W_{i}+\sum_{i=1}^{n}\left(\sum_{k=1}^{n}\epsilon^{\prime}_{\sigma_{k}(i)}\right)W^{\prime}_{i}\right)$
$=\epsilon\left(\frac{1}{n}\sum_{i=1}^{n}W_{i}\right)+(1-\epsilon)\left(\frac{1}{n}\sum_{i=1}^{n}W^{\prime}_{i}\right)$ (11)
where $\epsilon:=\sum_{j=1}^{n}\epsilon_{j}\in\{0,1\}$ satisfies $\mathbb{P}(\epsilon=1)=p$ and $\mathbb{P}(\epsilon=0)=1-p$.
Using (2), for any $t$ such that
$\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p)<t<\mathcal{H}(T_{\mu})(p),$
we have with Lemma 4:
$\mathbb{P}\left(\frac{1}{n}\sum_{k=1}^{n}X_{k}\geq t\right)$ $=p\,\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}W_{i}\geq t\right)+(1-p)\,\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}W_{i}^{\prime}\geq t\right)\underset{n\to\infty}{\longrightarrow}p.$
Similarly,
$\forall t<\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p):\ \ \mathbb{P}(Y_{n}\geq t)\to 1\qquad\operatorname{and}\qquad\forall t>\mathcal{H}(T_{\mu})(p):\ \ \mathbb{P}(Y_{n}\geq t)\to 0,$
which is exactly the profile of $S_{\mu,p}$. ∎
3 Classical sum concentration profiles
In practice, the tail quantile operator $T_{X}$ is rarely available in closed form. It is therefore natural to work instead with tractable envelopes for the survival operator $S_{\mu}$, and to propagate such bounds to $S_{X_{1}+\cdots+X_{n}}$ through Corollary 1.
For simplicity, for any $a>0$, let us introduce the operator $\mathrm{Id}^{-a}:\mathbb{R}\to 2^{\mathbb{R}}$ defined by
$\mathrm{Id}^{-a}(t)=\{t^{-a}\}\ \text{for }t>0,\qquad\mathrm{Id}^{-a}(t)=\emptyset\ \text{for }t\leq 0.$
Elementary integrations yield the following Hardy-transform identities:
$\forall q>1:\quad\mathcal{H}\bigl(\mathrm{Id}^{-\frac{1}{q}}\bigr)=\frac{q}{q-1}\,\mathrm{Id}^{-\frac{1}{q}}\qquad\text{and}\qquad\mathcal{H}(-\log)=1-\log.$ (12)
Inverting these identities (the order is preserved since all the mappings involved are nonincreasing), and applying Corollary 1, one obtains the following explicit profiles: for any constant $C>0$,
  • if $S_{\mu}\leq C\,\mathrm{Id}^{-q}$, then
    $S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq C\left(\frac{q}{q-1}\right)^{q}\mathrm{Id}^{-q};$
  • if $S_{\mu}\leq C\,\mathcal{E}_{1}$ where the notation $\mathcal{E}_{1}:t\mapsto\{e^{-t}\}$ was taken from Louart (2024), then
    $S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq C\,e\,\mathcal{E}_{1}.$ (13)
When regular inequalities like $S_{\mu}\leq C\,\mathrm{Id}^{-q}$ and $S_{\mu}\leq C\,\mathcal{E}_{1}$ are not satisfying or even reachable, it is worth trying to compare $S_{\mu}$ to $\mathrm{Id}^{-q}$ or $\mathcal{E}_{1}$ in the convex transformation order van Zwet (1964) to obtain simple and sharp bounds as shown in Corollary 1. We now prove this result.
Proof of Corollary 1.
Assume that $\mathrm{Id}^{-\frac{1}{q}}\circ\alpha$ is convex. Since $\alpha$ is nonincreasing and $\mathrm{Id}^{-\frac{1}{q}}$ is also nonincreasing, the composition $\mathrm{Id}^{-\frac{1}{q}}\circ\alpha$ is nondecreasing; therefore its inverse is concave. Noting that
$(\mathrm{Id}^{-\frac{1}{q}}\circ\alpha)^{-1}=\alpha^{-1}\circ\mathrm{Id}^{-q},$
we conclude that $\alpha^{-1}\circ\mathrm{Id}^{-q}$ is concave.
[…]
$\mathcal{H}(\alpha^{-1})(p)$ $=\int_{0}^{1}\alpha^{-1}(pr)\,dr=\int_{0}^{1}\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\bigl((pr)^{-\frac{1}{q}}\bigr)\,dr$
$\leq\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\!\left(\int_{0}^{1}(pr)^{-\frac{1}{q}}\,dr\right)=\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\bigl(\mathcal{H}(\mathrm{Id}^{-\frac{1}{q}})(p)\bigr),$
where the inequality is Jensen’s inequality applied to the concave mapping $\alpha^{-1}\circ\mathrm{Id}^{-q}$. Using (12), we get
$\mathcal{H}(T_{\mu})(p)\leq\mathcal{H}(\alpha^{-1})(p)\leq\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\!\left(\frac{q}{q-1}p^{-\frac{1}{q}}\right)=\alpha^{-1}\!\left(\left(\frac{q-1}{q}\right)^{q}p\right).$
Combining this with Theorem 1 in the i.i.d. marginal case gives
$T_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}(p)\leq\mathcal{H}(T_{\mu})(p)\leq\alpha^{-1}\!\left(\left(\frac{q-1}{q}\right)^{q}p\right).$
Inverting (and using again (1)) yields
$S_{X_{1}+\cdots+X_{n}}\leq\left(\frac{q}{q-1}\right)^{q}\,\alpha,$
which is the first result. The second result is shown the same way relying on the concavity of $\alpha^{-1}\circ\mathcal{E}_{1}$ and (13). ∎
Noting that
$\left(\frac{q}{q-1}\right)^{q}\underset{q\to\infty}{\longrightarrow}e,$
one may wonder whether the two convexity assumptions (power-type versus exponential-type) overlap. In general they do not; the next theorem shows that they coincide exactly in the limit $q\to\infty$.
Theorem 3.
Given a continuous mapping $f:(0,\infty)\to\mathbb{R}$, if for every $q>0$ the function $f\circ\mathrm{Id}^{-q}$ is convex (resp. concave), then the function $f\circ\mathcal{E}_{1}$ is convex (resp. concave). If we assume in addition that $f$ is nondecreasing (resp. nonincreasing), then the converse is true.
The theorem relies on the following limiting representation of the geometric mean.
Lemma 6 (Power–geometric limit).
Fix $a,b>0$ and $\lambda\in]0,1[$. Define, for $q>0$,
$m_{q,\lambda}(a,b):=\bigl(\lambda a^{-1/q}+(1-\lambda)b^{-1/q}\bigr)^{-q}.$
Then
$\lim_{q\to\infty}m_{q,\lambda}(a,b)=a^{\lambda}b^{1-\lambda}.$
Proof.
Set $r:=1/q$, $\alpha:=1/a$ and $\beta:=1/b$. Then
$m_{1/r,\lambda}(a,b)=\bigl(\lambda\alpha^{r}+(1-\lambda)\beta^{r}\bigr)^{-1/r},\qquad\text{so}\qquad-\log\bigl(m_{1/r,\lambda}(a,b)\bigr)=\frac{1}{r}\log\bigl(\lambda\alpha^{r}+(1-\lambda)\beta^{r}\bigr).$
Using the identity
$\frac{1}{r}\Bigl(\log g(r)-\log g(0)\Bigr)=\frac{1}{r}\int_{0}^{r}\frac{g^{\prime}(s)}{g(s)}\,ds,\qquad g(s):=\lambda\alpha^{s}+(1-\lambda)\beta^{s},$
(which appears for instance in Qi et al. (2000)), we obtain
$-\log\bigl(m_{1/r,\lambda}(a,b)\bigr)$ $=\frac{1}{r}\int_{0}^{r}\frac{\lambda\alpha^{s}\log(\alpha)+(1-\lambda)\beta^{s}\log(\beta)}{\lambda\alpha^{s}+(1-\lambda)\beta^{s}}\,ds$
$\underset{r\to 0}{\longrightarrow}\lambda\log(\alpha)+(1-\lambda)\log(\beta)=-\lambda\log(a)-(1-\lambda)\log(b),$
by continuity of the integrand. Taking exponentials yields the claim. ∎
Proof of Theorem 3.
We prove the statements for convexity; the concavity statements follow by applying the result to f-f.
Assume that for all $q>0$, the function $f\circ\mathrm{Id}^{-q}$ is convex. Fix $a,b>0$ and $\lambda\in[0,1]$. By convexity and by definition of $m_{q,\lambda}$,
$f\bigl(m_{q,\lambda}(a,b)\bigr)\leq\lambda f(a)+(1-\lambda)f(b).$
Letting $q\to\infty$ and using Lemma 6 together with continuity of $f$, we obtain
$f\bigl(a^{\lambda}b^{1-\lambda}\bigr)\leq\lambda f(a)+(1-\lambda)f(b).$
Writing $a=e^{-\alpha}$ and $b=e^{-\beta}$, this inequality is exactly the convexity of $f\circ\mathcal{E}_{1}$.
Conversely, assume that $f$ is nondecreasing and that $f\circ\mathcal{E}_{1}$ is convex. Fix $q>0$ and $a,b>0$, and write $a=e^{\alpha}$, $b=e^{\beta}$ with $\alpha=\log(a)$ and $\beta=\log(b)$. Since $\exp$ is convex, we have
$\lambda e^{\alpha}+(1-\lambda)e^{\beta}\geq e^{\lambda\alpha+(1-\lambda)\beta}.$
Raising to q-q (which reverses the inequality because $x\mapsto x^{-q}$ is nonincreasing), and using that $f$ is nondecreasing, we get
$f\bigl((\lambda a+(1-\lambda)b)^{-q}\bigr)$ $=f\bigl((\lambda e^{\alpha}+(1-\lambda)e^{\beta})^{-q}\bigr)$
$\leq f\bigl(e^{-q(\lambda\alpha+(1-\lambda)\beta)}\bigr).$
Finally, by convexity of $f\circ\mathcal{E}_{1}$,
$f\bigl(e^{-q(\lambda\alpha+(1-\lambda)\beta)}\bigr)$ $\leq\lambda f(e^{-q\alpha})+(1-\lambda)f(e^{-q\beta})=\lambda f(a^{-q})+(1-\lambda)f(b^{-q}),$
which proves that $f\circ\mathrm{Id}^{-q}$ is convex. ∎