### v2

# Let $X_{1},\ldots,X_{n}$ be random variables with an arbitrary dependence structure. Assume that there exists a nonincreasing mapping $\alpha:\mathbb{R}\to[0,1]$ with $\alpha(\mathbb{R})=[0,1]$ such that

[…]
1 Introduction and main results
Let $X_{1},\ldots,X_{n}$ be random variables with an arbitrary dependence structure. Assume that there exists a nonincreasing mapping $\alpha:\mathbb{R}\to[0,1]$ with $\alpha(\mathbb{R})=[0,1]$ such that
$\forall i\in[n],\ \forall t\in\mathbb{R}:\qquad\mathbb{P}(X_{i}\geq t)\leq\alpha(t).$
Then, as a direct consequence of a well-known property in the risk-measure literature, namely the subadditivity of expected shortfall, one obtains the universal bound (see Theorem 1.2)
$\mathbb{P}\!\left(\frac{1}{n}\sum_{i=1}^{n}X_{i}\geq t\right)\leq\bigl(\mathcal{H}(\alpha^{-1})\bigr)^{-1}(t),$
where $\alpha^{-1}$ denotes the generalized inverse of $\alpha$ and $\mathcal{H}(\alpha^{-1})$ is the Hardy transform, defined by
$\forall u\in(0,1]:\qquad\mathcal{H}(\alpha^{-1})(u):=\frac{1}{u}\int_{0}^{u}\alpha^{-1}(p)\,dp.$
The main contribution of this paper is to show, by an explicit construction, that without any additional assumption on the dependence structure the bounding profile $\bigl(\mathcal{H}(\alpha^{-1})\bigr)^{-1}$ is asymptotically optimal as $n\to\infty$ when $\alpha$ is the survival profile of the marginal law (see Theorems 1.82.1). More broadly, the operator formalism developed below covers nonincreasing, possibly discontinuous, survival profiles $\alpha$, and it naturally accommodates heterogeneous marginals, yielding a sharp universal analogue of the union bound for sums.
Following the approach of [Rockafellar and Royset(2014)] on superquantiles and our previous work on concentration inequalities [Louart(2024)], we encode concentration statements as inequalities between maximally nonincreasing operators rather than as pointwise inequalities between real-valued functions. The operator viewpoint is convenient here because objects underlying tail bounds (like survival functions and quantiles) are monotone but may have jumps (like the limiting profile of the law of large numbers); representing them as set-valued operators provides a canonical way to handle discontinuities without choosing arbitrary versions (lower or upper semicontinuous). In particular, atoms force a choice of left/right versions for survival functions and quantiles, and these choices can obscure inversion identities; interval-valued operators avoid this issue.
Concretely, we work in the class of “maximally nonincreasing operators” generalizing nonincreasing mappings on $\mathbb{R}$, which we denote $\mathcal{M}_{\downarrow}$ and which consists of set-valued mappings $\alpha:\mathbb{R}\to 2^{\mathbb{R}}$ satisfying [^2] The domain, range, and graph of an operator $\alpha:\mathbb{R}\to 2^{\mathbb{R}}$ are respectively defined as $\operatorname{Dom}(\alpha)=\{t\in\mathbb{R}\ |\ \alpha(t)\neq\emptyset\}$, $\operatorname{Ran}(\alpha)=\cup_{t\in\mathbb{R}}\alpha(t)$, and $\operatorname{Gra}(\alpha)=\{(x,y)\,:\,x\in\operatorname{Dom}(\alpha),\ y\in\alpha(x)\}$. The domain and range of a maximally nonincreasing operator are intervals; moreover, for any $t\in\mathbb{R}$, $\alpha(t)$ is a closed interval.:
$\forall(x,u),(y,v)\in\operatorname{Gra}(\alpha):\quad(y-x)(u-v)\geq 0\qquad\text{and}\qquad\operatorname{Dom}(\alpha)-\operatorname{Ran}(\alpha)=\mathbb{R}.$
Further characterizations can be found in [Bauschke and Combettes(2011), Louart(2024)].
In what follows, $\mu$ designates a probability measure on $\mathbb{R}$ and $X\sim\mu$ a generic random variable. Operator inversion allows one to move naturally between the survival operator $S_{X}:\mathbb{R}\to 2^{\mathbb{R}}$ (also denoted $S_{\mu}$) and the tail quantile operator $T_{X}:=S_{X}^{-1}:\mathbb{R}\to 2^{\mathbb{R}}$ (also denoted $T_{\mu}$), both maximally nonincreasing and defined respectively by
$\forall t\in\mathbb{R}:\qquad$ $S_{X}(t):=[\mathbb{P}(X>t),\mathbb{P}(X\geq t)]=[\mu(]t,+\infty[),\,\mu([t,+\infty[)]=:S_{\mu}(t)$
$\forall p\in[0,1]:\qquad$ $T_{X}(p):=\{t\in\mathbb{R}\ :\ p\in S_{X}(t)\}=:T_{\mu}(p).$
In particular, at an atom $t_{0}$, $S_{X}(t_{0})$ is the interval $\mathbb{P}(X>t_{0}),\mathbb{P}(X\geq t_{0})]$.
The natural order relation used to state concentration inequalities for survival operators was introduced in [Louart(2024)]:
Definition 1.1 (Interval Order and Point-wise resolvent Order between operators).
The order between intervals $A,B\subset\mathbb{R}$ is defined by [^3] For instance, if $A=[a_{1},a_{2}]$ and $B=[b_{1},b_{2}]$, then: $A\leq B\Leftrightarrow a_{1}\leq b_{1}\ \text{and}\ a_{2}\leq b_{2}$.
$A\leq B\quad\Longleftrightarrow\quad B_{+}\subset A_{+}\ \ \text{and}\ \ A_{-}\subset B_{-},$
with $A_{+}=\{x\in\mathbb{R}\ :\ \exists y\in A,\ y\leq x\}$ and $A_{-}=\{x\in\mathbb{R}\ :\ \exists y\in A,\ y\geq x\}$.
Given $f,g\in\mathcal{M}_{\downarrow}$, we write $f\leq g$ if and only if
$\operatorname{Dom}(f)\leq\operatorname{Dom}(g)\qquad\text{and}\qquad\forall x\in\operatorname{Dom}(f)\cap\operatorname{Dom}(g):\ f(x)\leq g(x),$
where $\operatorname{Dom}(f)$, $\operatorname{Dom}(g)$, $f(x)$ and $g(x)$ are (by construction) intervals (see [Bauschke and Combettes(2011), Proposition 20.31, Corollary 21.12]).
A convenient advantage of working with survival and tail quantile operators instead of working with the cumulative distribution operator $t\mapsto 1-S_{X}(t)$ and the quantile operator $p\mapsto T_{X}(1-p)$ as in [Rockafellar and Royset(2014)] is that an upper bound on the survival profile is equivalent to the corresponding upper bound on the tail quantile profile. More precisely, for any $\alpha\in\mathcal{M}_{\downarrow}$,
$S_{X}\leq\alpha\qquad\Longleftrightarrow\qquad T_{X}\leq\alpha^{-1}.$ (1)
Our objective can be stated as follows:
  • Task:
    Given $n$ probability measures $\mu_{1},\ldots,\mu_{n}$ on $\mathbb{R}$, construct a minimal [^4] For the pointwise resolvent order defined in Definition 1.1. operator $\alpha_{\mu_{1},\ldots,\mu_{n}}\in\mathcal{M}_{\downarrow}$ such that, for any collection of $n$ random variables $X_{1}\sim\mu_{1},\ldots,X_{n}\sim\mu_{n}$:
    $S_{X_{1}+\cdots+X_{n}}\leq\alpha_{\mu_{1},\ldots,\mu_{n}}.$
Assuming $\mu_{1}=\cdots=\mu_{n}=\mu$ and (i) considering $X_{1},\ldots,X_{n}$ i.i.d., (ii) considering $X_{1}=\cdots=X_{n}$ and (iii) using the union bound, one obtains the first simple bound [^5] Maximum of operators is given a rigorous definition in [Louart(2024)] that is not necessary to provide here since this inequality just plays a heuristic role.
$\max(\mathrm{Incr}_{\mathbb{E}[X]},S_{X})\leq\alpha_{\mu,\ldots,\mu}\leq nS_{X},$
where for any $\delta\in\mathbb{R}$, we denote by $\mathrm{Incr}_{\delta}\in\mathcal{M}_{\downarrow}$ the operator defined by
$\mathrm{Incr}_{\delta}(t)=\begin{cases}\{1\},&t<\delta,\\ [0,1],&t=\delta,\\ \{0\},&t>\delta.\end{cases}$
The naive upper bound $nS_{X}$ can of course be substantially improved. The optimal bound is expressed through the Hardy transform of the tail quantile operator $T_{X}$.
[…]
$\forall p\in\mathring{\operatorname{Dom}}(f)=\mathring{\operatorname{Dom}}(\mathcal{H}(f)):\qquad\mathcal{H}(f)(p)=\frac{1}{p}\int_{0}^{p}f(r)\,dr=\int_{0}^{1}f(pr)\,dr,$
then, since $\mathcal{H}(f)\in\mathcal{M}_{\downarrow}$, it can be shown that:
  • if $0\in\operatorname{Dom}(f)$, then $0\in\operatorname{Dom}(\mathcal{H}(f))$ and $\mathcal{H}(f)(0)=f(0)$,
  • denoting $p_{f}:=\sup(\operatorname{Dom}(f))$, we have $p_{f}\in\operatorname{Dom}(\mathcal{H}(f))$ and:
    $\mathcal{H}(f)(p_{f})=\left]-\infty,\frac{1}{p_{f}}\int_{0}^{p_{f}}f(r)\,dr\right]$
An example of a Hardy transform is depicted on Figure 1. We record a few elementary properties that will be useful later:
  • $0,p_{f}]\subset\operatorname{Dom}(\mathcal{H}(f))$ and for all $t\in]0,p_{f}[$, $\mathcal{H}(f)(t)$ is a singleton.
  • $\mathcal{H}$ is linear;
  • $f\leq\mathcal{H}(f)$;
  • if $f$ is convex then $\mathcal{H}(f)$ is convex;
  • if $f\leq g$ then $\mathcal{H}(f)\leq\mathcal{H}(g)$.
A first partial answer to our Task is, in fact, a reformulation of a property that is classical in financial mathematics: the subadditivity of expected shortfall [Acerbi and Tasche(2002)] (also called conditional value-at-risk [Rockafellar and Uryasev(2002)], and more recently superquantiles [Rockafellar and Royset(2014)]). Subadditivity is one of the coherence axioms introduced in [Artzner et al.(1999)Artzner, Delbaen, Eber, and Heath] for risk measures. This property was obtained simultaneously in [Acerbi and Tasche(2002), Proposition 3.1] (under the name “shortfall expectation”) and in [Rockafellar and Uryasev(2002), Corollary 12] (under the name “conditional value-at-risk”). Although these notions were initially presented differently, they were later shown to coincide; in our notation, they correspond to the mapping $p\mapsto\mathcal{H}(T_{X})(1-p)$. Subadditivity can be written as
$\mathcal{H}(T_{X_{1}+\cdots+X_{n}})\leq\mathcal{H}(T_{X_{1}})+\cdots+\mathcal{H}(T_{X_{n}}).$
Noting that $T_{X_{1}+\cdots+X_{n}}\leq\mathcal{H}(T_{X_{1}+\cdots+X_{n}})$ and relying on (1), we recover exactly:
Theorem 1.2.
Given $n$ random variables $X_{1},\ldots,X_{n}$ admitting expectations, one can bound:
$S_{X_{1}+\cdots+X_{n}}\leq\left(\mathcal{H}(T_{X_{1}})+\cdots+\mathcal{H}(T_{X_{n}})\right)^{-1}.$
As a comparison, in [Louart(2024), Proposition 42] we only set the weaker elementary bound (a generalization of the union bound to the non-identically distributed case):
$S_{X_{1}+\cdots+X_{n}}\leq n\bigl(T_{X_{1}}+\cdots+T_{X_{n}}\bigr)^{-1}.$
When $X_{1},\ldots,X_{n}\sim\mu$ are identically distributed, Theorem 1.2 rewrites:
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq\mathcal{H}(T_{\mu})^{-1}$ (2)
as expected, removing completely the dependence on $n$ of the bound.
Example 1.3 (Two-point distribution).
Let $X\sim\mu$ take only two values $a<b$, with
$\mathbb{P}(X=b)=\pi,\qquad\mathbb{P}(X=a)=1-\pi,\qquad\pi\in(0,1),$
so $\mathbb{E}[X]=a+\pi(b-a)$. Applying Theorem 1.2 yields the survival bound
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}(t)\leq\mathcal{H}(T_{\mu})^{-1}(t)=\begin{cases}\{1\},&t<\mathbb{E}[X],\\[4.0pt] \displaystyle\left\{\frac{\pi(b-a)}{t-a}\right\},&\mathbb{E}[X]<t<b,\\[8.0pt] [0,\pi],&t=b,\\[2.0pt] \{0\},&t>b.\end{cases}$
Remark 1.4.
While Theorem 1.2 is not designed to be tight at the level of moments, it is instructive to quantify what is lost when translating a tail bound into a moment bound. Still, it must be kept in mind that the strength of Theorem 1.2 is that it provides a universal tail bound, valid for all thresholds $t$ and all joint distributions of $(X_{1},\ldots,X_{n})$ with given marginals; for this reason, moments are not the most relevant metric here.
For any nonnegative random variable $X\sim\mu$, for some probability measure $\mu$ on $\mathbb{R}$ one has the identity (see for instance [Rockafellar and Royset(2014), (3.4)]):
$M_{q}(\mu):=\mathbb{E}[X^{q}]=\int_{0}^{\infty}\mathbb{P}(X^{q}\geq t)\,dt=\int_{0}^{1}T_{X}(p)^{q}\,dp.$ (3)
Following Hardy’s inequality [Hardy et al.(1952)Hardy, Littlewood, and Pólya, Theorem 9.8.2], valid for any measurable function $f:]0,1[\to\mathbb{R}_{+}$,
$\int_{0}^{1}\mathcal{H}(f)(t)^{q}\,dt\leq\left(\frac{q}{q-1}\right)^{q}\int_{0}^{1}f(t)^{q}\,dt.$
Given an operator $\alpha\in\mathcal{M}_{\downarrow}$ with $\operatorname{Dom}(\alpha)\subset\mathbb{R}_{+}$, we define the “moments” of $\alpha$ by
$M_{q}(\alpha):=\int_{\mathbb{R}_{+}}\alpha\!\left(t^{\frac{1}{q}}\right)\,dt,\quad q>0,$
so that $M_{q}(S_{X})=\mathbb{E}[X^{q}]=M_{q}(\mu)$.
Then, given $n$ random variables $X_{1},\ldots,X_{n}\sim\mu$, Theorem 1.2 applied in the $(X_{i})_{i\in[n]}$-identically distributed setting (2) yields the moment bound
$M_{q}\left(S_{\frac{1}{n}\sum_{i=1}^{n}X_{i}}\right)$ $=\int_{0}^{1}T_{\frac{1}{n}\sum_{i=1}^{n}X_{i}}(p)^{q}\,dp\leq\int_{0}^{1}\mathcal{H}(T_{X})(p)^{q}\,dp$
$\leq\left(\frac{q}{q-1}\right)^{q}\int_{0}^{1}T_{X}(p)^{q}\,dp=\left(\frac{q}{q-1}\right)^{q}M_{q}(S_{X}).$ (4)
This inequality is not tight for moments since a mere application of Jensen’s inequality already yields for all $q>1$ the moment optimal bound [^9] It is reached for instance in the case $X_{1}=\cdots=X_{n}$.:
$M_{q}(S_{\frac{1}{n}\sum_{i=1}^{n}X_{i}})\leq M_{q}(S_{X}).$
To pursue the study of the scope of Theorem 1.2, we give below a general result that efficiently bounds the concentration of the sum for a wide range of distributions. We employ the shorthand notation (already introduced in [Louart(2024)]) $\mathrm{Id}^{-a}:\mathbb{R}\to 2^{\mathbb{R}}$ defined by
$\forall t>0:\quad\mathrm{Id}^{-a}(t)=\{t^{-a}\},\qquad\text{and}\qquad\mathrm{Id}^{-a}((-\infty,0])=\emptyset.$ (5)
Corollary 1.5.
Given $n$ random variables $X_{1},\ldots,X_{n}$ and $\alpha\in\mathcal{M}_{\downarrow}$ satisfying $\operatorname{Ran}(\alpha)\subset\mathbb{R}^{+}_{*}$ and $q>1$ such that [^10] We say that an operator $g:\mathbb{R}\to 2^{\mathbb{R}}$ is convex if any real-valued mapping $h:\operatorname{Dom}(g)\to\mathbb{R}$ satisfying $\forall t\in\operatorname{Dom}(g)$, $h(t)\in g(t)$ is convex.:
$\forall i\in[n],\forall t\in\mathbb{R}:\quad\mathbb{P}(X_{i}\geq t)\leq\alpha(t)\qquad\text{and}\qquad\mathrm{Id}^{-\frac{1}{q}}\circ\alpha\ \ \text{is convex},$
one can bound
$\forall t\in\mathbb{R}:\quad\mathbb{P}\left(\frac{1}{n}\sum_{k=1}^{n}X_{k}\geq t\right)\leq\left(\frac{q}{q-1}\right)^{q}\alpha(t).$
If we assume in addition [^11] The second result of Proposition 3.2 in the next section states that $\log\circ\alpha$ convex implies $\mathrm{Id}^{-\frac{1}{q}}\circ\alpha$ convex for all $q>0$. that $\log\circ\alpha$ is convex, then
$\forall t\in\mathbb{R}:\quad\mathbb{P}\left(\frac{1}{n}\sum_{k=1}^{n}X_{k}\geq t\right)\leq e\alpha(t).$
This corollary is proved in Section 3.
We now turn to our main contribution: asymptotic sharpness of the inequality provided by Theorem 1.2. A general result for a triangular array of marginals $(\mu_{1}^{(n)},\ldots,\mu_{n}^{(n)})$ is stated in Section 2 (see Theorem 2.1); to keep the picture as simple as possible in this introduction, we first focus on the identically distributed case, where for each $n\in\mathbb{N}$ one has $\mu_{1}^{(n)}=\cdots=\mu_{n}^{(n)}=\mu$.
Before stating the result (under the assumption $\mathbb{E}[|X|]<\infty$), let us introduce the quantity
$\forall p\in]0,1[:\qquad\Delta_{\mu}(p):=\frac{\mathcal{H}(T_{\mu})(p)-\mathbb{E}[X]}{1-p}.$ (6)
Lemma 1.6.
Given any probability measure $\mu$ on $\mathbb{R}$ with $\mathbb{E}[|X|]<\infty$, any $X\sim\mu$, and any $p\in]0,1[$,
$\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p)\leq\mathbb{E}[X]\leq\mathcal{H}(T_{\mu})(p),$ (7)
with equality in (7) if and only if $\Delta_{\mu}(p)=0$.
Proof 1.7.
First note that $\Delta_{\mu}(p)\geq 0$ since $\mathcal{H}(T_{\mu})$ is nonincreasing and $\mathcal{H}(T_{\mu})(p)\geq\mathbb{E}[X]$ for $p\in(0,1)$. Moreover,
$\mathbb{E}[X]-\bigl(\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p)\bigr)=\frac{p}{1-p}\,\bigl(\mathcal{H}(T_{\mu})(p)-\mathbb{E}[X]\bigr)\geq 0,$
which proves (7) and the characterization of the equality case.
Given $p\in(0,1)$, define the limiting survival operator [^12] If $p=1$, $a_{\mu}(p)=b_{\mu}(p)=\mathbb{E}[X]$ and one can simply define $S_{\mu,1}:=\mathrm{Incr}_{\mathbb{E}[X]}$. $S_{\mu,p}\in\mathcal{M}_{\downarrow}$:
$S_{\mu,p}(t)=\begin{cases}\{1\},&\text{if }t<a_{\mu}(p),\\ [p,1],&\text{if }t=a_{\mu}(p),\\ \{p\},&\text{if }a_{\mu}(p)<t<b_{\mu}(p),\\ [0,p],&\text{if }t=b_{\mu}(p),\\ \{0\},&\text{if }t>b_{\mu}(p),\end{cases}$ $\text{with:}\quad\left\{\begin{aligned} a_{\mu}(p)&:=\mathcal{H}(T_{\mu})(p)-\Delta_{\mu}(p),\\ b_{\mu}(p)&:=\mathcal{H}(T_{\mu})(p).\end{aligned}\right$.
On the tail-quantile plot in Figure 1, the graph of $S_{\mu,p}^{-1}$ meets the graph of $\mathcal{H}(T_{\mu})$ at the point with abscissa $p$, shown as a green dot. The next theorem shows that such contact points can be achieved for any $p\in(0,1)$, so that the bound of Theorem 1.2 is asymptotically tight in this sense.
Figure 1: Representation of $T_{X}$, $\mathrm{Incr}_{\mathbb{E}[X]}$, $\mathcal{H}(T_{X})$ and $S_{\mu,p}^{-1}$ and the asymptotic contact point.
Theorem 1.8.
Given a probability distribution $\mu$ on $\mathbb{R}$ admitting a finite expectation, and given any $p\in(0,1]$, there exists for each $n\in\mathbb{N}$ a family of identically distributed random variables
$(X^{(n)}_{1},\ldots,X^{(n)}_{n})\qquad\text{with}\qquad X^{(n)}_{k}\sim\mu\ \text{for all }k\in[n],$
such that [^13] The set $\mathbb{R}\setminus\{a_{\mu}(p),b_{\mu}(p)\}$ is exactly the set of continuity points of $S_{\mu,p}$. for all $t\in\mathbb{R}\setminus\{a_{\mu}(p),b_{\mu}(p)\}$,
$S_{\frac{1}{n}\sum_{k=1}^{n}X^{(n)}_{k}}(t)\underset{n\to\infty}{\longrightarrow}S_{\mu,p}(t).$
This theorem is a direct consequence of Theorem 2.1 stated in Section 2.
The question remains open as to whether our sum concentration bound can be attained for finite $n$. One should presumably restrict to continuous distributions, since counterexamples are easy to find when $\mu$ has atoms. Moreover, when $n=2$, if we want to attain $S_{\mu,1}=\mathrm{Incr}_{\mathbb{E}[X]}$ with $S_{\frac{X_{1}+X_{2}}{2}}$ for $X_{1},X_{2}\sim\mu$, we must have $X_{1}=2\mathbb{E}[X]-X_{2}$ almost surely. Thus the concentration inequality can only be reached for $\mu$ symmetric around its expectation (and this is only for the threshold $p=1$). More freedom is available for $n\geq 3$, but the problem appears difficult and, having no specific application in mind, we do not explore this direction.
2 Sharpness for triangular array of marginals
In this section, we show that, under a uniform envelope assumption on marginal tails, the universal concentration bound from Theorem 1.2, applied for each $n$ to a triangular array of random variables $(X^{(n)}_{i})_{i\in[n]}$, is asymptotically sharp at every level $p\in]0,1[$. We will denote for simplicity
$\bar{X}^{(n)}=\frac{1}{n}\sum_{i=1}^{n}X_{i}^{(n)}.$
Moreover, given a family of $n$ measures $\nu=(\nu_{1},\ldots,\nu_{n})$, we denote the averaged tail-quantile operator
$T_{\nu}:=\frac{1}{n}\sum_{i=1}^{n}T_{\nu_{i}}\in\mathcal{M}_{\downarrow},$
and set
$b_{\nu}(p):=\mathcal{H}(T_{\nu})(p),\qquad\Delta_{\nu}(p):=\frac{b_{\nu}(p)-\int_{0}^{1}T_{\nu}(u)\,du}{1-p},\qquad a_{\nu}(p):=b_{\nu}(p)-\Delta_{\nu}(p).$
Theorem 2.1 (Asymptotic sharpness for non-identically distributed marginals).
Fix $p\in]0,1[$. For each $n\in\mathbb{N}$, let $\mu^{(n)}=(\mu^{(n)}_{1},\ldots,\mu^{(n)}_{n})$ be a family of probability measures on $\mathbb{R}$ such that there exists a measurable function $h:]0,1[\to\mathbb{R}_{+}$ such that
$\int_{0}^{1}h(u)\,du<\infty\qquad\text{and}\qquad\forall n\in\mathbb{N},\ \forall i\in[n],\ \forall u\in]0,1[:\quad T_{\mu_{i}^{(n)}}(u)\subset[-h(u),h(u)].$ (8)
Then, for each $n$, there exists a family $(X^{(n)}_{1},\ldots,X^{(n)}_{n})$ such that $X^{(n)}_{i}\sim\mu^{(n)}_{i}$ for all $i\in[n]$, and for any $\eta>0$ and any sequence $(t_{n})_{n\in\mathbb{N}}$ such that for all large $n$,
$t_{n}\leq a_{\mu^{(n)}}(p)-\eta,$ or $a_{\mu^{(n)}}(p)+\eta\leq t_{n}\leq b_{\mu^{(n)}}(p)-\eta,$ or $t_{n}\geq b_{\mu^{(n)}}(p)+\eta,$
one has respectively
$\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to 1,$ $\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to p,$ $\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to 0.$
Assumption (8) allows to use the dominated convergence theorem to set a law of large numbers in the proof of the theorem, it also ensures that $X^{(n)}_{i}$ for $n\in\mathbb{N}$, $i\in[n]$ all have finite expectations.
Proof 2.2.
Step 1: Quantile representation (push-forward form). For each $n\in\mathbb{N}$ and each $i\in[n]$, choose a measurable selection $\phi_{i}^{(n)}:[0,1]\to\mathbb{R}$ such that $\phi_{i}^{(n)}(u)\in T_{\mu_{i}^{(n)}}(u)$ for all $u\in[0,1]$. Equivalently, $\mu_{i}^{(n)}$ is the push-forward of the Lebesgue measure on $[0,1]$ through $\phi_{i}^{(n)}$.
Let $\epsilon$ be a Bernoulli random variable with $\mathbb{P}(\epsilon=1)=p$, and let $(U_{i})_{i\in[n]}$ be i.i.d. $\mathrm{Unif}(0,1)$, independent of $\epsilon$. Define, for all $i\in[n]$,
$U_{i}^{\epsilon}:=\epsilon\,(pU_{i})+(1-\epsilon)\,\bigl(p+(1-p)U_{i}\bigr),\qquad X_{i}^{(n)}:=\phi_{i}^{(n)}(U_{i}^{\epsilon}).$ (9)
By construction, conditional on $\epsilon$, the variables $(U_{i}^{\epsilon})_{i\in[n]}$ are i.i.d. and uniform on $[0,p]$ if $\epsilon=1$ and uniform on $[p,1]$ if $\epsilon=0$; moreover, unconditionally, each $U_{i}^{\epsilon}$ is $\mathrm{Unif}(0,1)$. Therefore, for each $i\in[n]$, $X_{i}^{(n)}\sim\mu_{i}^{(n)}$.
Step 2: Conditional expectations and identification of $a_{\mu^{(n)}}(p),b_{\mu^{(n)}}(p)$.
On $\{\epsilon=1\}$, $U_{i}^{\epsilon}=pU_{i}$, hence
$\mathbb{E}\bigl[X_{i}^{(n)}\mid\epsilon=1\bigr]=\int_{0}^{1}\phi_{i}^{(n)}(pu)\,du=\frac{1}{p}\int_{0}^{p}\phi_{i}^{(n)}(u)\,du=\mathcal{H}(T_{\mu_{i}^{(n)}})(p),$
where the last equality follows from the definition of the Aumann integral and the fact that $\mathcal{H}(T_{\mu_{i}^{(n)}})(p)$ is a singleton for $p\in]0,1[$. Averaging over $i$ and using linearity of the integral,
$\mathbb{E}\bigl[\bar{X}^{(n)}\mid\epsilon=1\bigr]$ $=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\bigl[X_{i}^{(n)}\mid\epsilon=1\bigr]$
$=\frac{1}{p}\int_{0}^{p}\left(\frac{1}{n}\sum_{i=1}^{n}T_{\mu_{i}^{(n)}}(u)\right)\,du=\mathcal{H}\!\left(T_{\mu^{(n)}}\right)(p)=b_{\mu^{(n)}}(p).$ (10)
Similarly, on $\{\epsilon=0\}$, $U_{i}^{\epsilon}=p+(1-p)U_{i}$, hence
$\mathbb{E}\bigl[X_{i}^{(n)}\mid\epsilon=0\bigr]=\int_{0}^{1}\phi_{i}^{(n)}\bigl(p+(1-p)u\bigr)\,du=\frac{1}{1-p}\int_{p}^{1}\phi_{i}^{(n)}(u)\,du,$
and summing over $i$ gives
$\mathbb{E}\bigl[\bar{X}^{(n)}\mid\epsilon=0\bigr]=\frac{1}{1-p}\int_{p}^{1}\left(\frac{1}{n}\sum_{i=1}^{n}T_{\mu_{i}^{(n)}}(u)\right)\,du=\frac{1}{1-p}\int_{p}^{1}T_{\mu^{(n)}}(u)\,du.$ (11)
On the other hand, by the identity $\mathbb{E}[Z]=\int_{0}^{1}T_{Z}(u)\,du$ for integrable $Z$ (see [Rockafellar and Royset(2014), (3.4)]),
$\mathbb{E}\bigl[\bar{X}^{(n)}\bigr]=\frac{1}{n}\sum_{i=1}^{n}\int_{0}^{1}T_{\mu_{i}^{(n)}}(u)\,du=\int_{0}^{1}T_{\mu^{(n)}}(u)\,du.$
Combining this with (2.2) and (11) yields
$\mathbb{E}\bigl[\bar{X}^{(n)}\mid\epsilon=0\bigr]=\frac{\mathbb{E}\bigl[\bar{X}^{(n)}\bigr]-p\,b_{\mu^{(n)}}(p)}{1-p}=b_{\mu^{(n)}}(p)-\frac{b_{\mu^{(n)}}(p)-\int_{0}^{1}T_{\mu^{(n)}}(u)\,du}{1-p}=a_{\mu^{(n)}}(p).$
Step 3: Conditional law of large numbers. We show that $\bar{X}^{(n)}\mid\{\epsilon=1\}\to b_{\mu^{(n)}}(p)$ in probability; the case $\epsilon=0$ is identical and yields $\bar{X}^{(n)}\mid\{\epsilon=0\}\to a_{\mu^{(n)}}(p)$.
Fix $t>0$ and work on the event $\{\epsilon=1\}$. Then $(X_{i}^{(n)})_{i\in[n]}$ are independent. For $M>0$, define the truncations
$X_{i}^{(n,M)}:=(-M)\vee X_{i}^{(n)}\wedge M,\qquad\bar{X}^{(n,M)}:=\frac{1}{n}\sum_{i=1}^{n}X_{i}^{(n,M)}.$
Fix $\delta>0$ and, without loss of generality, assume $\delta\leq 1$. Using a union bound, we obtain
$\mathbb{P}\Bigl(\bigl|\bar{X}^{(n)}-b_{\mu^{(n)}}(p)\bigr|\geq t\ \bigm|\ \epsilon=1\Bigr)$ $\leq\mathbb{P}\Bigl(\bigl|\bar{X}^{(n)}-\bar{X}^{(n,M)}\bigr|\geq t/3\ \bigm|\ \epsilon=1\Bigr)$ (12)
$\quad+\mathbb{P}\Bigl(\bigl|\bar{X}^{(n,M)}-\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]\bigr|\geq t/3\ \bigm|\ \epsilon=1\Bigr)$
$\quad+\mathbb{P}\Bigl(\bigl|\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]-b_{\mu^{(n)}}(p)\bigr|\geq t/3\Bigr).$
We choose $M$ large enough so that the first term is at most $\delta/2$ (uniformly in $n$) and the third term is equal to 0; then we let $n\to\infty$ to control the second term via Hoeffding’s inequality.
We start with the bound
$\bigl|\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]-\mathbb{E}[\bar{X}^{(n)}\mid\epsilon=1]\bigr|$ $\leq\mathbb{E}\bigl[|\bar{X}^{(n)}-\bar{X}^{(n,M)}|\mid\epsilon=1\bigr]$
$\leq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\bigl[|X_{i}^{(n)}-X_{i}^{(n,M)}|\mid\epsilon=1\bigr],$ (13)
and $\mathbb{E}[\bar{X}^{(n)}\mid\epsilon=1]=b_{\mu^{(n)}}(p)$ by (2.2). Hence the third term in (12) is zero as soon as
$\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\bigl[|X_{i}^{(n)}-X_{i}^{(n,M)}|\mid\epsilon=1\bigr]<t/3.$
Moreover, since $|X_{i}^{(n)}-X_{i}^{(n,M)}|\leq|X_{i}^{(n)}|\,\mathbf{1}_{\{|X_{i}^{(n)}|>M\}}$, using the quantile representation on $\{\epsilon=1\}$ we have
$\mathbb{E}\bigl[|X_{i}^{(n)}-X_{i}^{(n,M)}|\mid\epsilon=1\bigr]\leq\frac{1}{p}\int_{0}^{p}|\phi_{i}^{(n)}(u)|\,\mathbf{1}_{\{|\phi_{i}^{(n)}(u)|>M\}}\,du\leq\frac{1}{p}\int_{0}^{p}h(u)\,\mathbf{1}_{\{h(u)>M\}}\,du.$
The right-hand side tends to 0 as $M\to\infty$ by dominated convergence (since $\int_{0}^{1}h(u)\,du<\infty$). Thus, for the given $t>0$ and $\delta\in(0,1]$, one can choose $M$ such that
$\frac{1}{p}\int_{0}^{p}h(u)\,\mathbf{1}_{\{h(u)>M\}}\,du\leq\frac{\delta t}{6},$
which implies both
$\mathbb{P}\Bigl(\bigl|\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]-b_{\mu^{(n)}}(p)\bigr|\geq t/3\Bigr)=0,$
and, by Markov’s inequality and (2.2),
$\mathbb{P}\Bigl(\bigl|\bar{X}^{(n)}-\bar{X}^{(n,M)}\bigr|\geq t/3\ \bigm|\ \epsilon=1\Bigr)\leq\frac{3}{t}\,\mathbb{E}\bigl[|\bar{X}^{(n)}-\bar{X}^{(n,M)}|\mid\epsilon=1\bigr]\leq\frac{\delta}{2},$
uniformly in $n$.
Finally, on $\{\epsilon=1\}$ the truncated variables satisfy $X_{i}^{(n,M)}\in[-M,M]$ and remain independent; thus Hoeffding’s inequality gives
$\mathbb{P}\Bigl(\bigl|\bar{X}^{(n,M)}-\mathbb{E}[\bar{X}^{(n,M)}\mid\epsilon=1]\bigr|\geq t/3\ \bigm|\ \epsilon=1\Bigr)\leq 2\exp\!\Bigl(-\frac{nt^{2}}{18M^{2}}\Bigr)\leq\frac{\delta}{2},$ (14)
for $n$ large enough. Plugging these bounds into (12) yields
$\mathbb{P}\Bigl(\bigl|\bar{X}^{(n)}-b_{\mu^{(n)}}(p)\bigr|\geq t\ \bigm|\ \epsilon=1\Bigr)\leq\delta,$
which proves $\bar{X}^{(n)}\mid\{\epsilon=1\}\to b_{\mu^{(n)}}(p)$ in probability. The case $\epsilon=0$ is identical and yields $\bar{X}^{(n)}\mid\{\epsilon=0\}\to a_{\mu^{(n)}}(p)$.
Step 4: Tail profile away from the discontinuity points. Let $\eta>0$ and $(t_{n})_{n}$ be as in the statement.
If $t_{n}\leq a_{\mu^{(n)}}(p)-\eta$, then on $\{\epsilon=0\}$ we have $\bar{X}^{(n)}\to a_{\mu^{(n)}}(p)$ in probability and therefore $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=0)\to 1$; similarly on $\{\epsilon=1\}$ since $b_{\mu^{(n)}}(p)\geq a_{\mu^{(n)}}(p)$, we also have $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=1)\to 1$. Hence $\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to 1$.
If $t_{n}\geq b_{\mu^{(n)}}(p)+\eta$, then $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=1)\to 0$ and $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=0)\to 0$, hence $\mathbb{P}(\bar{X}^{(n)}\geq t_{n})\to 0$.
Finally, if $a_{\mu^{(n)}}(p)+\eta\leq t_{n}\leq b_{\mu^{(n)}}(p)-\eta$, then $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=1)\to 1$ and $\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=0)\to 0$, hence
$\mathbb{P}(\bar{X}^{(n)}\geq t_{n})=p\,\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=1)+(1-p)\,\mathbb{P}(\bar{X}^{(n)}\geq t_{n}\mid\epsilon=0)\to p.$
This proves the three limits announced.
3 Classical sum concentration profiles
In practice, the tail quantile operator $T_{X}$ is rarely available in closed form. It is therefore natural to work instead with tractable envelopes for the survival operator $S_{\mu}$, and to propagate such bounds to $S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}$ through Corollary 1.5.
Elementary calculations yield the following Hardy-transform identities (with the power-operator notation introduced in (5)):
$\forall q>1:\quad\mathcal{H}\bigl(\mathrm{Id}^{-\frac{1}{q}}\bigr)=\frac{q}{q-1}\,\mathrm{Id}^{-\frac{1}{q}}\qquad\text{and}\qquad\mathcal{H}(-\log)=1-\log.$ (15)
Inverting these identities (inversion preserves the order since all the mappings involved are nonincreasing), and applying Corollary 1.5, one obtains the following explicit profiles: for any constant $C>0$,
  • if $S_{\mu}\leq C\,\mathrm{Id}^{-q}$, then
    $S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq C\left(\frac{q}{q-1}\right)^{q}\mathrm{Id}^{-q};$
  • if $S_{\mu}\leq C\,\mathcal{E}_{1}$, where the notation $\mathcal{E}_{1}:t\mapsto\{e^{-t}\}$ is taken from [Louart(2024)], then
    $S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq C\,e\,\mathcal{E}_{1}.$ (16)
When simple pointwise bounds such as $S_{\mu}\leq C\,\mathrm{Id}^{-q}$ or $S_{\mu}\leq C\,\mathcal{E}_{1}$ are not satisfactory, or are not available, it is worth comparing $S_{\mu}$ to $\mathrm{Id}^{-q}$ or $\mathcal{E}_{1}$ with respect to the convex transformation order [van Zwet(1964)]. This yields simple and sharp bounds, as stated in Corollary 1.5, which we now prove.
Proof 3.1 (Proof of Corollary 1.5).
Assume that $\mathrm{Id}^{-\frac{1}{q}}\circ\alpha$ is convex. Since $\alpha$ is nonincreasing and $\mathrm{Id}^{-\frac{1}{q}}$ is also nonincreasing, the composition $\mathrm{Id}^{-\frac{1}{q}}\circ\alpha$ is nondecreasing; therefore its inverse is concave. Noting that
$(\mathrm{Id}^{-\frac{1}{q}}\circ\alpha)^{-1}=\alpha^{-1}\circ\mathrm{Id}^{-q},$
we conclude that $\alpha^{-1}\circ\mathrm{Id}^{-q}$ is concave.
Now assume that for all $i\in[n]$$S_{X_{i}}\leq\alpha$. By (1), this implies $T_{X_{i}}\leq\alpha^{-1}$, and since the Hardy transform is order-preserving, we obtain $\mathcal{H}(T_{X_{i}})\leq\mathcal{H}(\alpha^{-1})$. For $p>0$,
$\mathcal{H}(\alpha^{-1})(p)$ $=\int_{0}^{1}\alpha^{-1}(pr)\,dr=\int_{0}^{1}\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\bigl((pr)^{-\frac{1}{q}}\bigr)\,dr$
$\leq\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\!\left(\int_{0}^{1}(pr)^{-\frac{1}{q}}\,dr\right)=\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\bigl(\mathcal{H}(\mathrm{Id}^{-\frac{1}{q}})(p)\bigr),$
where the inequality is Jensen’s inequality applied to the concave mapping $\alpha^{-1}\circ\mathrm{Id}^{-q}$. Using (15), we get
$\mathcal{H}\left(\frac{1}{n}\sum_{k=1}^{n}T_{X_{k}}\right)(p)\leq\mathcal{H}(\alpha^{-1})(p)\leq\bigl(\alpha^{-1}\circ\mathrm{Id}^{-q}\bigr)\!\left(\frac{q}{q-1}p^{-\frac{1}{q}}\right)=\alpha^{-1}\left(\left(\frac{q-1}{q}\right)^{q}p\right).$
Combining this with Theorem 1.2 in the identically distributed case gives
$T_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}(p)\leq\mathcal{H}\left(\frac{1}{n}\sum_{k=1}^{n}T_{X_{k}}\right)(p)\leq\alpha^{-1}\!\left(\left(\frac{q-1}{q}\right)^{q}p\right).$
Inverting (and using again (1)) yields
$S_{\frac{1}{n}\sum_{k=1}^{n}X_{k}}\leq\left(\frac{q}{q-1}\right)^{q}\,\alpha,$
which is the first result.
The second result is proved analogously, using the concavity of $\alpha^{-1}\circ\mathcal{E}_{1}$ (equivalently, the convexity of $\log\circ\alpha$) and the identity $\mathcal{H}(-\log)=1-\log$ from (15).
Noting that
$\left(\frac{q}{q-1}\right)^{q}\underset{q\to\infty}{\longrightarrow}e,$
one may wonder whether the two convexity assumptions (power-type versus exponential-type) overlap. In general they do not; the next proposition shows that they coincide exactly in the limit $q\to\infty$.
Proposition 3.2.
Given a continuous function $f:(0,\infty)\to\mathbb{R}$, if for every $q>0$ the function $f\circ\mathrm{Id}^{-q}$ is convex (resp. concave), then the function $f\circ\mathcal{E}_{1}$ is convex (resp. concave). If we assume in addition that $f$ is nondecreasing (resp. nonincreasing), then the converse is true.
The proposition relies on the following limiting representation of the geometric mean.
Lemma 3.3 (Power–geometric limit).
Fix $a,b>0$ and $\lambda\in]0,1[$. Define, for $q>0$,
$m_{q,\lambda}(a,b):=\bigl(\lambda a^{-1/q}+(1-\lambda)b^{-1/q}\bigr)^{-q}.$
Then
$\lim_{q\to\infty}m_{q,\lambda}(a,b)=a^{\lambda}b^{1-\lambda}.$
Proof 3.4.
Set $r:=1/q$, $A:=1/a$ and $B:=1/b$. Then
$m_{1/r,\lambda}(a,b)=\bigl(\lambda A^{r}+(1-\lambda)B^{r}\bigr)^{-1/r},\qquad\text{so}\qquad-\log\bigl(m_{1/r,\lambda}(a,b)\bigr)=\frac{1}{r}\log\bigl(\lambda A^{r}+(1-\lambda)B^{r}\bigr).$
Using the identity
$\frac{1}{r}\Bigl(\log g(r)-\log g(0)\Bigr)=\frac{1}{r}\int_{0}^{r}\frac{g^{\prime}(s)}{g(s)}\,ds,\qquad g(s):=\lambda A^{s}+(1-\lambda)B^{s},$
(which appears for instance in [Qi et al.(2000)Qi, Mei, Xia, and Xu]), we obtain
$-\log\bigl(m_{1/r,\lambda}(a,b)\bigr)$ $=\frac{1}{r}\int_{0}^{r}\frac{\lambda A^{s}\log(A)+(1-\lambda)B^{s}\log(B)}{\lambda A^{s}+(1-\lambda)B^{s}}\,ds$
$\underset{r\to 0}{\longrightarrow}\lambda\log(A)+(1-\lambda)\log(B)=-\lambda\log(a)-(1-\lambda)\log(b),$
by continuity of the integrand. Taking exponentials yields the claim.
Proof 3.5 (Proof of Proposition 3.2).
We prove the statements for convexity; the concavity statements follow by applying the result to f-f.
Assume that for all $q>0$, the function $f\circ\mathrm{Id}^{-q}$ is convex. Fix $a,b>0$ and $\lambda\in[0,1]$. By convexity and by definition of $m_{q,\lambda}$,
$f\bigl(m_{q,\lambda}(a,b)\bigr)\leq\lambda f(a)+(1-\lambda)f(b).$
Letting $q\to\infty$ and using Lemma 3.3 together with continuity of $f$, we obtain
$f\bigl(a^{\lambda}b^{1-\lambda}\bigr)\leq\lambda f(a)+(1-\lambda)f(b).$
Writing $a=e^{-x}=\mathcal{E}_{1}(x)$ and $b=\mathcal{E}_{1}(y)$, $a^{\lambda}b^{1-\lambda}=\mathcal{E}_{1}(\lambda x+(1-\lambda)y)$ and this inequality is exactly the convexity of $f\circ\mathcal{E}_{1}$.
Conversely, assume that $f$ is nondecreasing and that $f\circ\mathcal{E}_{1}$ is convex. Fix $q>0$ and $a,b>0$, and write $a=e^{x}$, $b=e^{y}$ with $x=\log(a)$ and $y=\log(b)$. Since $\exp$ is convex, we have
$\lambda e^{x}+(1-\lambda)e^{y}\geq e^{\lambda x+(1-\lambda)y}.$
Raising to q-q (which reverses the inequality because $x\mapsto x^{-q}$ is nonincreasing), and using that $f$ is nondecreasing, we get
$f\bigl((\lambda a+(1-\lambda)b)^{-q}\bigr)$ $=f\bigl((\lambda e^{x}+(1-\lambda)e^{y})^{-q}\bigr)$
$\leq f\bigl(e^{-q(\lambda x+(1-\lambda)y)}\bigr).$
Finally, by convexity of $f\circ\mathcal{E}_{1}$,
$f\bigl(e^{-q(\lambda x+(1-\lambda)y)}\bigr)$ $\leq\lambda f(e^{-qx})+(1-\lambda)f(e^{-qy})=\lambda f(a^{-q})+(1-\lambda)f(b^{-q}),$
which proves that $f\circ\mathrm{Id}^{-q}$ is convex.